OpenStudy (anonymous):

Given that n≥1, what is the smallest positive integer n whose divisors have the product of n^8?

myininaya (myininaya):

\[(1)^4(1)^4=(1)^8\] 1 is the smallest integer n i don't think i understand what you are asking

OpenStudy (anonymous):

it has to be greater then one because n>1

OpenStudy (anonymous):

It says n >= 1

myininaya (myininaya):

it has = 1

OpenStudy (anonymous):

n>1 so it cannt equal one

OpenStudy (anonymous):

So, you are saying you typed the question wrong? Because your original typed question says n>=1.

myininaya (myininaya):

\[n \ge 1 => \text{ n can be =1 or greater than 1}\]

OpenStudy (anonymous):

Given that n>1, what is the smallest positive integer n whose divisors have a produst of n^8?